One of two small classrooms is chosen at random with equally likely probability, and then a student is chosen at random from the chosen classroom. Classroom #1 has 5 boys and 13 girls. Classroom #2 has 15 boys and 9 girls. What is the probability that Classroom #2 was chosen at random, given that a girl was chosen? Your answers should be rounded to 4 digits after the decimal.

A 0.6582

B 0.3418

C 0.1875

D 0.3750

E 0.2708

F None of the above
Answer #1

P(girl) = P(girl | classroom 1) * P(classroom 1) + P(girl | classroom 2) * P(classroom 2)

          = (13/18) * (1/2) + (9/24) * (1/2)

          = 0.5486

P(classroom 2 | girl) = P(girl | classroom 2) * P(classroom 2) / P(girl) = (9/24) * (1/2) / 0.5486 = 0.3418

Option-B) 0.3418